OpenStudy (sayan):

What is the answer? |3x-2|<=|2x-5|

OpenStudy (anonymous):

Sayan, if you graph the two you will get two 'v' shapes, and you're looking for the area in which the |3x-2| line is below the |2x-5| line. You can find where they intersect to find the bounds on x for which there are solutions

myininaya (myininaya):

Remember that \[|x|=\sqrt{x^2}\] so we have \[\sqrt{(3x-2)^2} \le \sqrt{(2x-5)^2}\] squaring both sides gives: \[(3x-2)^2 \le (2x-5)^2\] \[9x^2-12x+4 \le 4x^2-20x+25\] \[5x^2+8x-21 \le 0\] \[5x^2+15x-7x-21 \le 0\] \[5x(x+3)-7(x+3) \le 0\] \[(x+3)(5x-7) \le 0\] so =0 when x=-3 and x=7/5 now i like to to make a number line and test where the inequality is true ----------|---------------------|-------------------------- -3 7/5 so choose a number before -3 like -4 plug in -4 to see if inequality holds now plug in 0 to see if inequality holds (note you chould have chose any number btw -3 and 7/5) now last interval: choose a number afterr 7/5 and see if the inequality holds

myininaya (myininaya):

|3(-4)-2|<|2(-4)-5| does not hold so (-inf,-3) is not a part of our solution |3(0)-2|<|2(0)-5| does hold so (-3,7/5) is a part of our solution |3(3)-2|<|2(3)-5| does not hold so (7/5, inf) is not a part our solution so the solution is the set: (-3,7/5)